OpenStudy (anonymous):

solve by factoring please show the steps on getting the answer. 1. 2x^2+x=10

OpenStudy (anthonyym):

First set it equal to 0 to put it in standard form ax^2 + bx + c

OpenStudy (anonymous):

factor please help 1.a^3-125= 2.2x^2+X-4= 3.3X^2+X-4= 4.X^4negative2X^2-3=

OpenStudy (anthonyym):

So which question do you need help on? Want to do the one you posted first?